# Bilinear MLPs enable weight-based mechanistic interpretability

### v1

@@ 4.4 Adversarial masks: general attacks from weights
We construct an adversarial mask from an eigenvector for the digit 3 (Figure 7A). Even though the original eigenvector resembles the digit, the pseudoinverse-based mask does not (see Appendix K for more examples). The accuracy, averaged over masks from the top three eigenvectors, drops significantly more than the baseline of randomly permuting the mask, despite regularizing the model during training using dense Gaussian noise with a standard deviation of 0.15. The corresponding rise in misclassification indicates effective steering towards the adversarial digit.
@@ 5 Language: finding interactions between SAE features
Each output of a bilinear layer is described by weighted pairwise interactions between their input features. Previous sections show that this can be successfully leveraged to trace between a bilinear layer’s inputs and outputs. Here, we turn towards tracing between latent feature dictionaries obtained by training sparse autoencoders (SAEs) on the MLP inputs or outputs for a 6-layer bilinear transformer trained on TinyStories (Eldan & Li, 2023) (see training details in Appendix F).
@@ 5.1 Sentiment negation circuit
The sentiment negation circuit computes the activation of two opposing output features in layer 4 (index 1882 and 1179) that form a fully linear subspace. The cosine similarity of their decoder vectors is -0.975. Based on their top activations, the output features activate on negation tokens (“not”, “never”, “wasn’t”) and boosts either positive sentiment tokens (“good”, “safe”, “nice”) or negative sentiment tokens (“bad”, “hurt”, “sad”), so we denote the two features as the not-good and the not-bad features respectively. See Appendix M for the top activations of all features mentioned.
Figure 8: The sentiment negation circuit for computing the not-good and not-bad output features. A) The interaction submatrix containing the top 15 interactions. B) The top interacting features projected onto the top positive and negative eigenvectors using cosine similarity, with the symbols for different clusters matching the labels in A. Clusters coincide with the projection of meaningful directions such as the difference in “bad” vs “good” token unembeddings and the MLP’s input activations for input tokens “[BOS] not”. C) The not-good feature activation compared to its approximation by the top two eigenvectors.
Focusing on the not-good feature, the top interactions for computing its activations resemble an AND-gate (Figure 8A). Input features that boost negative sentiment tokens (blue squares) have strong positive interactions with negation token features (green triangles), but both have negligible self-interactions. So both types of input features are needed to activate the not-good feature and flip the boosting from negative to positive sentiment. The one positive sentiment feature (orange downward triangle) has interactions with the opposite sign. The interactions shown are significantly larger than the typical cross-interactions with a standard deviation of 0.004 (Figure 24.
@@
Figure 9: Activation correlations with low-rank approximations. A) Average correlation over output features computed over all inputs before the SAE’s non-linearity (TopK) is applied or only on inputs where the feature is active. B) The distribution of active-only correlations for approximations using the top two eigenvectors. C) Scatter plots for a random set of nine output features. Approximations use the top two eigenvectors.
@@ 5.2 Low-rank approximations of output feature activations
The top eigenvectors can be used to approximate the activations of the SAE output features using a truncated form of Equation 3. To focus on the more meaningful tail of large activations, we compute the approximation’s correlation conditioned on the output feature being active. For comparison, we also compute the approximation’s correlation with the output feature activations before the SAE’s nonlinearity (TopK) is applied. These correlations are generally lower because they are dominated by inputs with low activations.
We find that features are surprisingly low-rank, with the average active-only correlation starting at 0.67 for approximations by a single eigenvector and rising steadily with additional eigenvectors (Figure 9A). Most features (69%) have a high correlation (>0.75>0.75) even when approximated by just two eigenvectors (Figure 9B). Scatter plots for a random sample of features show that the low-rank approximation often captures the tail dependence well (Figure 9C). Overall these results suggest that the interactions that produce large output activations are low-rank, making their interpretability potentially easier.
@@ 6 Discussion
Summary. This paper introduces a novel approach to weight-based interpretability that leverages the close-to-linear structure of bilinear layers. A key result is that, using an eigenvector decomposition, we can identify the most important input directions that explain the layer’s output along a given direction. The top eigenvectors are often interpretable, for example for MNIST they function as edge-detectors for strokes specific to each digit. The lack of element-wise nonlinearity in bilinear MLPs allows us to transform their weights into interaction matrices that connect input to output features and then extract the low-rank structure. In language models, we find that many SAE output features are well-approximated by low-rank interaction matrices, particularly at large activations. We highlighted one example of an extracted low-rank circuit that flips the sentiment of the next token if the current token is a negation (“not”). The behavior of this circuit can be easily understood in terms of the top eigenvectors, whereas finding a similar circuit in conventional MLPs would be more difficult. Overall, our results demonstrate that bilinear MLPs offer intrinsic interpretability that can aid in feature and circuit extraction.
Implications. The main implication of our work is that weight-based interpretability is viable, even for large language models. Bilinear MLPs can replace conventional MLPs in transformers with minimal cost while offering intrinsic interpretability due to their lack of element-wise nonlinearities and close-to-linear structure. Current circuit analysis techniques rely on gradient-based approximations (Syed et al., 2023; Marks et al., 2024) or use transcoders (Dunefsky et al., 2024) to approximate MLPs. Both approaches depend on an input dataset, potentially leading to poor performance out-of-distribution, and they may not fully capture the nonlinear computations in MLPs. In contrast, bilinear MLPs can be transformed into explicit feature interaction matrices and decomposed in a way fully equivalent to the original computations. Extracting interactions more directly from the weights should lead to better, more robust circuits. Weight-based interpretability may also offer better safety guarantees since we could plausibly prove bounds on a layer’s outputs based by quantifying the residual weights not captured in a circuit’s interactions.
@@ Acknowledgements
We are grateful to Narmeen Oozeer, Nora Belrose, Philippe Chlenski, and Kola Ayonrinde for useful feedback on the draft. We are grateful to the AI Safety Camp program where this work first started and to the ML Alignment & Theory Scholars (MATS) program that supported Michael and Alice while working on this project. We thank CoreWeave for providing compute for the finetuning experiments. This research received funding from the Flemish Government under the ”Onderzoeksprogramma Artificiële Intelligentie (AI) Vlaanderen” programme.
@@
This paper proposes no advancements to the state-of-the-art in model capabilities. Rather, it provides new methods to analyze the internals of models to increase our understanding. The only potential misuse the authors envision is using this technique to leak details about the dataset that the model has learned more efficiently. However, this can be avoided by using this technique during safety evaluation.
@@
We aspired to make this work as reproducible as possible. First, Appendix F (among others) aims to provide detailed and sufficient descriptions to independently recreate our training setups. Second, our code (currently public but not referenced for anonymity) contains separate files that can be used to generate the figures in this paper independently. We ensured setting seeds across training runs such that recreated figures would be equivalent. Third, all models that are compute-intensive to train, such as the SAEs and the LMs, will be shared publicly. Lastly, we will publish an interactive demo, which will allow independent analysis of the figures in Appendix A, Appendix B, and Appendix M in a way this document cannot.
@@ B.1 Regularization
Weight decay generally acts as a sparsifier for eigenvalues but does not significantly impact the eigenvectors. This is extremely useful as it can zero out the long tail of unimportant eigenvalues, strongly reducing the labor required to analyze a model fully (more details in Appendix D).
Dropout did not seem to impact our models. Overfitting was still an issue, even for very high values (>0.5>0.5). We suspect this may change in larger or capacity-constrained models.
@@ Appendix C Explainability: a small case study with eigenvectors
Appendix D Sparsity: weight decay versus input noise
@@
Appendix E Truncation & similarity: a comparison across sizes
@@
Appendix F Experimental setups: a detailed description
@@
F.1 Image classification setup
@@
F.2 Language model setup
The language model used in this paper is a 6-layer modern transformer model (Touvron et al., 2023) where the SwiGLU is replaced with a bilinear MLP (Figure 20). The model has about 33 million parameters. The training setup is detailed in Table 2. As the training dataset, we use a simplified and cleaned version of TinyStories (Eldan & Li, 2023) that remedies the following issues.
@@ F.2 Language model setup
Table 2: Model architecture and training setup. Omitted parameters are the HuggingFace defaults.
Transformer Training Parameters
@@
F.3 Sparse autoencoder setup
All discussed SAEs use a TopK activation function, as described in Gao et al. (2024). We found $k=30$ to strike a good balance between sparseness and reconstruction loss. This paper studies quite narrow dictionaries (4x expansion) for simplicity. The exact hyperparameters are shown in Table 3, and the attained loss added (Equation 4) across layers is shown in Figure 21.
@@ F.3 Sparse autoencoder setup
Appendix G Bilinear transformers: a loss comparison
While experiments on large models show bilinear layers to only marginally lag behind SwiGLU (Shazeer, 2020), this section quantifies this accuracy trade-off through compute efficiency. We performed our experiments using a 6-layer transformer model trained on TinyStories. For these experiments, we use $\text{d\_model}=512$ and $\text{d\_hidden}=2048$, resulting in roughly 30 million parameters. However, we have found these results to hold across all sizes we tried.
@@ Appendix G Bilinear transformers: a loss comparison
Appendix H Finetuning: your transformer is secretly bilinear
@@
Appendix I Tensor decompositions: extracting shared features
I.1 Re-expressing the bilinear tensor
Given a complete or over-complete set of $m$ $\bm{u}$}-vectors, we can re-express $\bm{\mathsfit{B}}$} in terms of the eigenvectors, which amounts to a change of basis. To avoid multiple contributions from similar $\bm{u}$}-vectors, we have to use the pseudo-inverse, which generalizes the inverse for non-square matrices. Taking the $\bm{u}$}-vectors as the columns of $\bm{U}$}, the pseudo-inverse $\bm{U}$$^{+}$ satisfies ${\bm{U}}{\bm{U}}^{+}=I$, as long as $\bm{U}$} is full rank (equal to $d$). Then
${\bm{\mathsfit{B}}}$ $=\sum^{m}_{k}{\bm{u}}^{+}_{:k}\otimes{\bm{Q}}_{k}$ (5)
@@ I.1 Re-expressing the bilinear tensor
where ${\bm{u}}^{+}_{:k}$ are the rows of $U^{+}$ and $Q_{k}=\sum_{i}\lambda_{\{k,i\}}{\bm{v}}_{\{k,i\}}\otimes{\bm{v}}_{\{k,i\}}$ is the eigendecomposition of the interaction matrix corresponding for ${\bm{u}}{k:}$. We can then recover the interaction matrices from ${\bm{Q}}_{k}={\bm{u}}_{k:}\cdot_{\text{out}}{\bm{\mathsfit{B}}}$ using the fact that ${\bm{u}}_{k:}\cdot{\bm{u}}^{+}_{:k^{\prime}}=\delta_{kk^{\prime}}$ (Kronecker delta). Note that the eigenvectors within a single output direction $k$ are orthogonal but will overlap when comparing across different output directions.
Appendix J Bilinear layers: a practical guide
@@
Appendix K Adversarial masks: additional figure
@@
Appendix L Language: further details for feature circuits
@@
Appendix M Input features of the negation feature

### v2

@@ 4.4 Adversarial masks: general attacks from weights
We construct an adversarial mask from an eigenvector for the digit 3 (Figure 7A). Even though the original eigenvector resembles the digit, the pseudoinverse-based mask does not (see Appendix M for more examples). The accuracy, averaged over masks from the top three eigenvectors, drops significantly more than the baseline of randomly permuting the mask despite regularizing the model during training using dense Gaussian noise with a standard deviation of 0.15. The corresponding rise in misclassification indicates effective steering towards the adversarial digit.
@@ 5 Language: finding interactions between SAE features
Each output of a bilinear layer is described by weighted pairwise interactions between their input features. Previous sections show that this can be successfully leveraged to trace between a bilinear layer’s inputs and outputs. Here, we turn towards tracing between latent feature dictionaries obtained by training sparse autoencoders (SAEs) on the MLP inputs or outputs for a 6-layer bilinear transformer trained on TinyStories (Eldan & Li, 2023) (see training details in Appendix G).
@@ 5.1 Sentiment negation circuit
The sentiment negation circuit computes the activation of two opposing output features in layer 4 (index 1882 and 1179) that form a fully linear subspace. The cosine similarity of their decoder vectors is -0.975. Based on their top activations, the output features activate on negation tokens (“not”, “never”, “wasn’t”) and boosts either positive sentiment tokens (“good”, “safe”, “nice”) or negative sentiment tokens (“bad”, “hurt”, “sad”), so we denote the two features as the not-good and the not-bad features respectively. See Appendix O for the top activations of all features mentioned.
Figure 8: The sentiment negation circuit that computes the not-good and not-bad output features. A) The interaction submatrix containing the top 15 interactions. B) The projection of top interacting features onto the top eigenvectors using cosine similarity. The symbols for different clusters match the labels in A. Clusters coincide with the projection of meaningful directions such as the difference in “bad” vs “good” token unembeddings and the MLP’s input activations for the input “[BOS] not”. C) The not-good feature activation compared to its approximation by the top two eigenvectors.
Focusing on the not-good feature, the top interactions for computing its activations resemble an AND-gate (Figure 8A). Input features that boost negative sentiment tokens (blue squares) have strong positive interactions with negation token features (green triangles), but both have negligible self-interactions. So, both types of input features are needed to activate the not-good feature and flip the boost from negative to positive sentiment. The one positive sentiment feature (orange downward triangle) interacts with the opposite sign. The interactions shown are significantly larger than the typical cross-interactions with a standard deviation of 0.004 (Figure 27.
@@ 5.2 Low-rank approximations of output feature activations
The top eigenvectors can be used to approximate the activations of the SAE output features using a truncated form of Equation 3. To focus on the more meaningful tail of large activations, we compute the approximation’s correlation conditioned on the output feature being active. The correlations of inactive features are generally lower because they are dominated by ‘noise’. We evaluate this on three bilinear transformers at approximately 2/3 depth: a 6-layer TinyStories (‘ts-tiny’) and two FineWeb models with 12 and 16 layers (‘fw-small’ and ‘fw-medium’).
We find that features are surprisingly low-rank, with the average correlation starting around 0.65 for approximations by a single eigenvector and rising steadily with additional eigenvectors (Figure 9A). Most features have a high correlation ($>0.75$) even when approximated by just two eigenvectors (Figure 9B). Scatter plots for a random sample of features show that the low-rank approximation often captures the tail dependence well (Figure 9C). Interestingly, we find the approximation to drastically improve with longer SAE training times while other metrics change only slightly. This indicates a ‘hidden’ transition near convergence and is further discussed in Appendix H. Overall, these results suggest that the interactions that produce large output activations are low-rank, making their interpretability potentially easier.
[…]
Figure 9: Activation correlations with low-rank approximations for differently-sized transformers. A) Average correlation over output features computed over every input where the feature is active. B) The distribution of active-only correlations for approximations using the top two eigenvectors. C) Scatter plots for a random set of nine output features on ‘fw-medium’. Approximations use the top two eigenvectors. Low correlation scores generally only occur on low-activation features.
@@ 6 Discussion
Summary. This paper introduces a novel approach to weight-based interpretability that leverages the close-to-linear structure of bilinear layers. A key result is that we can identify the most important input directions that explain the layer’s output along a given direction using an eigenvector decomposition. The top eigenvectors are often interpretable, for example for MNIST they function as edge-detectors for strokes specific to each digit. The lack of element-wise nonlinearity in bilinear MLPs allows us to transform their weights into interaction matrices that connect input to output features and then extract the low-rank structure. In language models, we find that many SAE output features are well-approximated by low-rank interaction matrices, particularly at large activations. We highlighted one example of an extracted low-rank circuit that flips the sentiment of the next token if the current token is a negation (“not”). The behavior of this circuit can be easily understood in terms of the top eigenvectors, whereas finding a similar circuit in conventional MLPs would be more difficult. Overall, our results demonstrate that bilinear MLPs offer intrinsic interpretability that can aid in feature and circuit extraction.
Implications. The main implication of our work is that weight-based interpretability is viable, even for large language models. Bilinear MLPs can replace conventional MLPs in transformers with minimal cost while offering intrinsic interpretability due to their lack of element-wise nonlinearities and close-to-linear structure. Current circuit analysis techniques rely on gradient-based approximations (Syed et al., 2023; Marks et al., 2024) or use transcoders (Dunefsky et al., 2024) to approximate MLPs. Both approaches depend on an input dataset, potentially leading to poor performance out-of-distribution, and they may not fully capture the nonlinear computations in MLPs. In contrast, bilinear MLPs can be transformed into explicit feature interaction matrices and decomposed in a way fully equivalent to the original computations. Extracting interactions more directly from the weights should lead to better, more robust circuits. Weight-based interpretability may also offer better safety guarantees since we could plausibly prove bounds on a layer’s outputs by quantifying the residual weights not captured in a circuit’s interactions.
@@ Acknowledgements
We are grateful to Narmeen Oozeer, Nora Belrose, Philippe Chlenski, and Kola Ayonrinde for helpful feedback on the draft. We are grateful to the AI Safety Camp program where this work first started and to the ML Alignment & Theory Scholars (MATS) program that supported Michael and Alice while working on this project. We thank CoreWeave for providing compute for the finetuning experiments. This research received funding from the Flemish Government under the ”Onderzoeksprogramma Artificiële Intelligentie (AI) Vlaanderen” programme.
Contributions
Michael performed the bulk of the work on the MNIST analysis and provided valuable insights across all presented topics. Thomas worked on the Language Models section and was responsible for code infrastructure. The paper was written in tandem, each focusing on their respective section.
@@
This paper proposes no advancements to the state-of-the-art in model capabilities. Rather, it provides new methods to analyze the internals of models to increase our understanding. The only misuse the authors envision is using this technique to leak details about the dataset that the model has learned more efficiently. However, this can be avoided by using this technique during safety evaluation.
@@
We aspired to make this work as reproducible as possible. First, Appendix G (among others) aims to provide detailed and sufficient descriptions to independently recreate our training setups. Second, our code (currently public but not referenced for anonymity) contains separate files that can be used to generate the figures in this paper independently. We used seeds across training runs so that recreated figures would be equivalent. Third, all models that are compute-intensive to train, such as the SAEs and the LMs, will be shared publicly. Lastly, we will publish an interactive demo, which will allow independent analysis of the figures in Appendix A, Appendix B, and Appendix O in a way this document cannot.
@@ B.1 Regularization
Weight decay generally acts as a sparsifier for eigenvalues but does not significantly impact the eigenvectors. This is extremely useful as it can zero out the long tail of unimportant eigenvalues, strongly reducing the labor required to analyze a model fully (more details in Appendix E).
Dropout did not seem to impact our models. Overfitting was still an issue, even for very high values ($>0.5$). We suspect this may change in larger or capacity-constrained models.
@@ Appendix C Explainability: a small case study with eigenvectors
Appendix D HOSVD: finding the most important shared features
[…]
Instead of contributing to a single output dimension, each interaction matrix can contribute to an arbitrary direction, shown at the bottom right (”contributions”). Further, the importance of the contributions is determined by their singular value, which is shown at the top right. The remainder of the visualization shows the top eigenvalues and the corresponding eigenvectors.
The most important output direction separates digits with a prominent vertical line (1, 4, and 7) from digits with a prominent horizontal line (5 specifically). Similarly, the second most important direction splits horizontal from vertical lines but is more localized to the top half. Specifically, it splits by the orientation of the top stroke (whether it starts/ends left or right).
Figure 17: The most important output direction of an MNIST model roughly splits digits by its horizontality or verticality.
Figure 18: The second most important output direction of an MNIST model splits digits according to the orientation of its top stroke.
We observe that the output directions uncovered through HOSVD somewhat correspond to meaningful concepts, albeit sometimes dominated by a specific digit (such as 5 and 6). Less significant directions often highlight specific portions of digits that seem meaningful but are more challenging to describe. In summary, while in the case of MNIST, the results are not particularly more interpretable than decomposing according to digits, we believe this technique increases in utility (but also computational cost) as the number of output classes increases.
Appendix E Sparsity: weight decay versus input noise
@@
Appendix F Truncation & similarity: a comparison across sizes
@@
Appendix G Experimental setups: a detailed description
@@
G.1 Image classification setup
@@
G.2 Language model setup
The language model used in section 5 is a 6-layer modern transformer model (Touvron et al., 2023) where the SwiGLU is replaced with a bilinear MLP (Figure 22). The model has about 33 million parameters. The training setup is detailed in Table 3. As the training dataset, we use a simplified and cleaned version of TinyStories (Eldan & Li, 2023) that remedies the following issues.
@@ G.2 Language model setup
Table 2: Tinystories training setup. Omitted parameters are the HuggingFace defaults.
TinyStories Training Parameters
@@
The models used in the experiments shown in Figure 9 are trained of the FineWeb dataset (Penedo et al., 2024). These follow the architecture of GPT2-small (12 layers) and GPT2-medium (16 layers) but have bilinear MLPs. Their parameter count is 162M and 335M, respectively. Both use the Mixtral tokenizer.
Table 3: FineWeb training setup. Omitted parameters are the HuggingFace defaults.
FineWeb Training Parameters
weight decay 0.1
batch size 512
context length 512
learning rate 6e-4
optimizer AdamW
schedule linear decay
tokens $\pm$ 32B
initialisation gpt2
G.3 Sparse autoencoder setup
All discussed SAEs use a TopK activation function, as described in Gao et al. (2024). We found $k=30$ to strike a good balance between sparseness and reconstruction loss. section 5 studies quite narrow dictionaries (4x expansion) for simplicity. The exact hyperparameters are shown in Table 4, and the attained loss added (Equation 4) across layers is shown in Figure 23.
@@ G.3 Sparse autoencoder setup
Appendix H Correlation: analyzing the impact of training time
Figure 9 shows how a few eigenvectors capture the essence of SAE features. This section discusses the impact of SAE quality, measured through training steps, on the resulting correlation. In short, we find features of SAEs that are trained longer are better approximated with few eigenvectors.
We train 5 SAEs with an expansion factor of 16 on the output of the MLP at layer 12 of the ‘fw-medium’ model. Each is trained twice as long as the last. The feature approximation correlations are computed over 100K activations; features with less than 10 activations (of which there are less than 1000) are considered dead and not shown. The reconstruction error and loss recovered between SAEs differ only by 10% while the correlation mean changes drastically (Table 5). The correlation distribution is strongly bimodal for the ‘under-trained’ SAEs (shown in Figure 24 with darker colors). Given more training time, this distribution shifts towards higher correlations. The activation frequencies of features are mostly uncorrelated with their approximations.
Table 5: The SAE metrics along with the mean of the correlations shown in Figure 24. The correlation improves strongly with longer training, while other metrics only change marginally.
Training steps (relative) 1 2 4 8 16
Normalized MSE 0.17 0.16 0.16 0.15 0.15
Loss recovered 0.60 0.61 0.65 0.65 0.66
Correlation mean 0.17 0.28 0.42 0.52 0.59
Figure 24: Feature approximation correlations using two eigenvectors across SAEs with different training times. Darker is shorter, and bright is longer. This shows a clear bimodal distribution for ‘under-trained’ SAEs, which vanishes upon longer training, indicating some form of convergence.
Appendix I Bilinear transformers: a loss comparison
[…]
@@ Appendix I Bilinear transformers: a loss comparison
Appendix J Finetuning: your transformer is secretly bilinear
@@
Appendix K Tensor decompositions: extracting shared features
[…]
@@
[…]
Appendix L Bilinear layers: a practical guide
@@
Appendix M Adversarial masks: additional figure
@@
Appendix N Language: further details for feature circuits
@@
Appendix O Input features of the negation feature